OpenStudy (anonymous):

-7square root-2x+17=-7

OpenStudy (anonymous):

\[-7\sqrt{-2x+17}=-7 \] Is that the equation In which you are trying to solve? X would have to equal 8. Devide both sides by -7. \[\sqrt{17-2x}\]=1 Now square both sides. 17-2x=1 Isolate the variable -2x=-16 Divide by -2 x=16

OpenStudy (anonymous):

just a typo on the last line... but it's correct, btw he mentioned it already prior to his solution...